4 years ago marion was 2/3 as old as Les was. Now she is 3/4 as old as he is. How old is each now
Answer by Maths68(1474)

4 years ago marion was 2/3 as old as Les was. Now she is 3/4 as old as he is. How old is each now
Let
Present age of Marion = m
Present age of Les = l


Marion is 3/4 as old as Les is
m=(3/4)(l)
m=3l/4.................................(1)

4 years ago
Marion was = m-4
Les was = l-4
Then
(m-4)=2/3(l-4)
3(m-4)=2(l-4)
3m-12=2l-8
3m=2l-8+12
3m=2l+4................................(2)
Put the value of m from (1) to (2)
3m=2l+4
3(3l/4)=2l+4
9l/4=2l+4
9l=4(2l+4)
9l=8l+16
9l-8l=16
l=16
Put the value of l in (1)
m=3l/4.................................(1)
m=(3)(16)/4
m=(3)(4)
m=12


Present age of Marion = m = 12 years old
Present age of Les = l = 16 years old